# Scan-Adaptive Dynamic MRI Undersampling Using a Dictionary of Efficiently Learned Patterns

### v1

@@ 1 Introduction
Early MRI acceleration relied on pulse sequence design [21, 5, 39] and parallel imaging (p-MRI) [29, 44], which uses multi-coil sensitivity encoding but can suffer from noise amplification at high acceleration. Compressed sensing (CS) addresses this regime by reconstructing from undersampled measurements using sparsity-based regularization [9, 24, 23]. The earliest work for dynamic MRI reconstruction used the temporal redundancy across frames, motivating methods based on temporal transform sparsity, such as - BLAST and - SENSE [38, 17], - ISD [20], and - SPARSE-SENSE [10]. Later works incorporated a low rank prior into the reconstruction framework that included - SLR [22] and L+S [26]. More recent dynamic MRI methods exploit manifold/low-rank structure in -space, including methods like SToRM [27], KLR [25], bilinear manifold [37], and LTSA [8].
@@
In accelerated MRI, acquisition design plays a critical role in determining image quality. Standard compressed sensing (CS) techniques typically rely on fixed heuristic patterns, such as variable-density random sampling (VDRS) [24] or equispaced sampling [14], to promote incoherent aliasing. Moving beyond these fixed designs, early optimization techniques derived sampling patterns from static dataset statistics, utilizing metrics such as power spectra and experimental design criteria [19, 32, 15, 36].
@@
On the other hand, scan-adaptive techniques such as SeqMRI [43], MNet [16], SUNO [11] tailor acquisition to the individual subject. For instance, SeqMRI formulates sampling as a sequential decision process, optimizing a policy to select -space lines iteratively based on the current reconstruction state. Similarly, MNet uses a CNN to map low-frequency -space data directly to a scan-specific Cartesian mask. While effective for static imaging, these methods face significant barriers in dynamic cardiac MRI. The high dimensionality of dynamic -space renders sequential decision-making and direct mask prediction computationally prohibitive or unstable for ensuring spatiotemporal consistency across cardiac phases.
To address this, we propose Dynamic Scan-Adaptive MRI Undersampling using Neighbor-based Optimization (dSUNO), which extends our prior SUNO framework [11] to dynamic cardiac MRI. The proposed dSUNO framework replaces the original ICD-based mask optimization with a novel Randomized Batched Iterative Coordinate Descent (RB-ICD) algorithm that updates subsets of phase-encoding lines simultaneously instead of one line at time. This batched updating reduces optimization time, making scan-adaptive mask learning practical for cardiac cine series. Second, dSUNO extends the original test-time neighbor selection from static 2D images to a spatiotemporal neighbor search over entire cardiac cine series, enabling scan-adaptive mask estimation while maintaining temporal consistency across frames.
@@ 3.1.2 OCMR Dataset
The OCMR dataset [6] contains multi-coil cine cardiac MRI data acquired on Siemens platforms (Prisma 3T, Avanto 1.5T, and Sola 1.5T), including both fully sampled and prospectively undersampled scans. We focused on the short-axis (sax) cine scans from the fully sampled subset acquired at 1.5T using an 18-channel coil. Each sequence contains 20–30 temporal frames.
For this dataset, the data was split into 139 training and 26 testing slices. Retrospective undersampling masks were applied to simulate acceleration factors of 4 and 6. This dataset served as a benchmark for evaluating undersampling and reconstruction strategies.
@@ 3.3 Reconstruction Network Training
We trained both MoDL and MostNet models on the MCardiac and OCMR datasets using the optimized scan-adaptive ICD masks at 4, 8, and 12 acceleration. Each dynamic scan was split into training and validation sets.
For MoDL [1], we used a UNet [33] as the denoiser inside the training framework, in which unrollings of the denoiser and the conjugate gradient (CG) blocks were used. The regularization parameter controlling the weighting of the two terms in MoDL was set to , and the tolerance for the CG algorithm used was . For MostNet, we used six cascades of unrolled architecture using the CRNN-based spatiotemporal regularization. The relative weighting of the CRNN prior and the regularization parameter was set to and , respectively. The network was trained using an loss and the Adam optimizer with a learning rate of for 100 epochs. Sensitivity analyses justifying the choice of the regularization parameter and the number of unrolled iterations are provided in the Supplementary document.
@@ 6 Conclusion
This work presented a scan-adaptive Cartesian undersampling framework for dynamic cardiac MRI, in which sampling masks are optimized offline using an alternating sampler–reconstructor training technique and selected at test time via a neighbor search. Mask optimization is performed using a novel randomized batched iterative coordinate descent (RB-ICD) algorithm. The proposed MostNet architecture integrates data-consistency layers from MoDL with temporal regularization modules inspired by CTFNet.
Experimental results on both the public OCMR dataset and the in-house MCardiac dataset show that the proposed scan-adaptive sampling technique consistently improves reconstruction quality over commonly used Cartesian sampling baselines across multiple acceleration factors. Compared to existing baseline sampling patterns, the proposed technique reduces aliasing artifacts and better preserves temporal consistency across frames. Radiologist evaluations further indicate that scan-adaptive masks lead to improved diagnostic image quality under aggressive undersampling.
The results suggest that tailoring Cartesian sampling patterns to individual dynamic cardiac MRI scans can provide benefits over fixed sampling patterns. Future work will investigate extensions to non-Cartesian and physics-informed sampling trajectories, as well as closer integration with self-supervised and deep image prior–based reconstruction methods.
@@ Acknowledgment
The authors acknowledge Dr. Maryam Sayadi and Evan Bell for helpful discussions.

### v2

@@ 1 Introduction
Early MRI acceleration relied on pulse sequence design [32, 5, 53] and parallel imaging (p-MRI) [41, 58, 10], which uses multi-coil sensitivity encoding but can suffer from noise amplification at high acceleration. Compressed sensing (CS) addresses this regime by reconstructing from undersampled measurements using sparsity-based regularization [13, 14, 35, 34]. The earliest work for dynamic MRI reconstruction used the temporal redundancy across frames, motivating methods based on temporal transform sparsity, such as - BLAST and - SENSE [52, 26], - ISD [31], and - SPARSE-SENSE [16]. Later works incorporated a low rank prior into the reconstruction framework that included - SLR [33] and L+S [38]. More recent dynamic MRI methods exploit manifold/low-rank structure in -space, including methods like SToRM [39], KLR [37], bilinear manifold [50], and LTSA [12].
@@
Recent advances have explored implicit neural networks that map spatial and temporal coordinates to image intensities for cardiac MRI reconstruction. Earlier works such as NeRP [49], neural implicit k-space [23, 24], and Fourier-feature-based models [29] show how implicit representations can reconstruct dynamic cardiac motion directly from sparse measurements. Subsequent methods introduced subspace constraints [24], spatiotemporal coordinate encoding [15], and self-supervised or physics-informed training [51, 7] to improve reconstruction quality and generalization further.
In accelerated MRI, acquisition design plays a critical role in determining image quality. Standard compressed sensing (CS) techniques typically rely on fixed heuristic patterns such as variable-density random sampling (VDRS) [35], equispaced sampling [21], Poisson-disc sampling [6], and hybrid variable-density Poisson-disc masks [36, 30] to promote incoherent aliasing. Moving beyond these fixed designs, early optimization techniques derived sampling patterns from static dataset statistics, utilizing metrics such as power spectra and experimental design criteria [28, 44, 22, 48].
@@
On the other hand, scan-adaptive techniques such as SeqMRI [57], MNet [25], SUNO [17, 11, 18] tailor acquisition to the individual subject. For instance, SeqMRI formulates sampling as a sequential decision process, optimizing a policy to select -space lines iteratively based on the current reconstruction state. Similarly, MNet uses a CNN to map low-frequency -space data directly to a scan-specific Cartesian mask. While effective for static imaging, these methods face significant barriers in dynamic cardiac MRI. The high dimensionality of dynamic -space renders sequential decision-making and direct mask prediction computationally prohibitive or unstable for ensuring spatiotemporal consistency across cardiac phases.
To address this, we propose Dynamic Scan-Adaptive MRI Undersampling using Neighbor-based Optimization (dSUNO), which extends our prior SUNO framework [18] to dynamic cardiac MRI. The proposed dSUNO framework replaces the original ICD-based mask optimization with a novel Randomized Batched Iterative Coordinate Descent (RB-ICD) algorithm that updates subsets of phase-encoding lines simultaneously instead of one line at a time. This batched updating reduces optimization time, making scan-adaptive mask learning practical for cardiac cine series. Second, dSUNO extends the original test-time neighbor selection from static 2D images to a spatiotemporal neighbor search over entire cardiac cine series, enabling scan-adaptive mask estimation while maintaining temporal consistency across frames.
@@ 3.1.2 OCMR Dataset
The OCMR dataset [8] contains multi-coil cine cardiac MRI data acquired on Siemens platforms (Prisma 3T, Avanto 1.5T, and Sola 1.5T), including both fully sampled and prospectively undersampled scans. We focused on the short-axis (sax) cine scans from the fully sampled subset acquired at 1.5T using an 18-channel coil. Each sequence contains 20–30 temporal frames. For this dataset, the data was split into 139 training and 26 testing slices. Retrospective undersampling masks were applied to simulate acceleration factors of 4 and 6. This dataset served as a benchmark for evaluating undersampling and reconstruction strategies.
@@ 3.3 Reconstruction Network Training
We trained both MoDL and MostNet models on the MCardiac and OCMR datasets using the optimized scan-adaptive ICD masks at 4, 8, and 12 acceleration. Each dynamic scan was split into training and validation sets. For MoDL [1], we used a UNet [45] as the denoiser inside the training framework, in which six unrollings of the denoiser and the conjugate gradient (CG) blocks were used. The regularization parameter controlling the weighting of the two terms in MoDL was set to , and the tolerance for the CG algorithm used was . For MostNet, we used six cascades of unrolled architecture using the CRNN-based spatiotemporal regularization. The relative weighting of the CRNN prior and the regularization parameter was set to and , respectively. The network was trained using an loss and the Adam optimizer with a learning rate of for 100 epochs. Sensitivity analyses justifying the choice of the regularization parameter and the number of unrolled iterations are provided in the Supplementary document.
@@ 6 Conclusion
This work presented a scan-adaptive Cartesian undersampling framework for dynamic cardiac MRI, in which sampling masks are optimized offline using an alternating sampler–reconstructor training technique and selected at test time via a neighbor search. Mask optimization is performed using a novel randomized batched iterative coordinate descent (RB-ICD) algorithm. The proposed MostNet architecture integrates data-consistency layers from MoDL with temporal regularization modules inspired by CTFNet. Experimental results on both the public OCMR dataset and the in-house MCardiac dataset show that the proposed scan-adaptive sampling technique consistently improves reconstruction quality over commonly used Cartesian sampling baselines across multiple acceleration factors. Compared to existing baseline sampling patterns, the proposed technique reduces aliasing artifacts and better preserves temporal consistency across frames. Radiologist evaluations further indicate that scan-adaptive masks lead to improved diagnostic image quality under aggressive undersampling. The results suggest that tailoring Cartesian sampling patterns to individual dynamic cardiac MRI scans can provide benefits over fixed sampling patterns. Future work will investigate extensions to non-Cartesian and physics-informed sampling trajectories, as well as closer integration with self-supervised and deep image prior–based reconstruction methods.
@@ Acknowledgment
The authors gratefully acknowledge Prof. Maryam Sayadi (Michigan State University) for helpful discussions. The authors also acknowledge Evan Bell (currently at Johns Hopkins University) for contributions to this work while he was at Michigan State University.